Find the solution to the system by the addition (elimination)method
16x-6y=-28 (1)
4x+12y=14 (2)

Answer by SanDiegoMath(2)

(1) 16x-6y = -28
(2) 4x+12y = 14
Dividing (2)'s coefficients in half:
(1) 16x-6y = -28
(3) 2x+6y = 7
Adding (1) and (3) cancels the 'y' term:
(4) 18x = -21
Dividing by 18 we get the value of x:
(5) x = -21/18 = -7/6
Substituting into (3):
(6) 2(-7/6) + 6y = 7
or -7/3 + 6y = 7
solving for y:
6y = 7 + 7/3 = (21/3) + (7/3) = 28/3
y = 28/18 = 14/9
Solution: x=-7/6 and y =14/9
Verifying:
(1) 16x-6y = 16(-7/6) - 6(14/9) = -(56/3)-(28/3) = -84/3 = -28 , it works
(2) 4x+12y = 4(-7/6) + 12(14/9) = -(14/3) + (56/3) = (56-14)/3 = 42/3 = 14 , it works, too.